The probability of a man hitting a target is 0.40. If he fires 5 times then, find the probability of hitting the target, (a) exactly 3 times, (b) at least 4 times


1
Expert's answer
2021-04-29T17:20:46-0400

p=0.4,q=1-p=0.6

n=5


(a) P(exactly three times)= "^5C_3(0.4)^3(0.6)2=10\times 0.4^3\times 0.36=0.2304"


(b) P( at least 4 times)="P(r\ge 4)=P(r=4)+P(r=5)=^5C_4(0.4)^4(0.6)+^5C_5(0.4)^5"


"=3\times (0.4)^4+(0.4)5=0.08704"